# Customize-A-Video: One-Shot Motion Customization of Text-to-Video Diffusion Models

### v1

@@ Abstract
Image customization has been extensively studied in text-to-image (T2I) diffusion models, leading to impressive outcomes and applications. With the emergence of text-to-video (T2V) diffusion models, its temporal counterpart, motion customization, has not yet been well investigated. To address the challenge of one-shot motion customization, we propose Customize-A-Video that models the motion from a single reference video and adapting it to new subjects and scenes with both spatial and temporal varieties. It leverages low-rank adaptation (LoRA) on temporal attention layers to tailor the pre-trained T2V diffusion model for specific motion modeling from the reference videos. To disentangle the spatial and temporal information during the training pipeline, we introduce a novel concept of appearance absorbers that detach the original appearance from the single reference video prior to motion learning. Our proposed method can be easily extended to various downstream tasks, including custom video generation and editing, video appearance customization, and multiple motion combination, in a plug-and-play fashion. Our project page can be found at https://anonymous-314.github.io.
Figure 1: Customize-A-Video takes as input a single reference video (top left) and transfers its motion onto new generated videos with plausible variance. (1-2) Results of transferring the dancing twirling from the lady onto Ironman with two random output variants. (3) Results of transferring the motion onto multiple subjects. (4) Results of combining multiple motion customization together, i.e., both dancing twirling and with aerial camera flight over. (5) Results of combining proposed motion customization and existing image customization methods to support both appearance and motion customization.
Figure 2: Our proposed Temporal LoRA and its training and inference processes. All noise and denoising schedules are omitted for simplicity. (1) We bypass all temporal layers in a base T2V diffusion model and apply appearance absorber such as S-LoRA or Textual Inversion on its spatial attention layers. The module is trained on unordered video frames. (2) We apply T-LoRA on all temporal attentions in the full base T2V model. The trained appearance absorber is also loaded and frozen. The module is trained on the target video data. (3) During inference, only the trained T-LoRA is loaded. A new video with the customized motion is generated by a prompt describing the new appearance and the desired motion.
@@ 1 Introduction
Replicating an iconic motion in novel scenes is highly desirable for video creation. Recent large-scale diffusion-based text-to-video (T2V) generation models [43, 5] demonstrate impressive outcomes in generating imaginative videos based on text depictions. However, they struggle with precise motion control and often demand extensive prompt engineering. Another thread of work on video editing such as [48, 9, 45, 4] leverages large generative models’ generalization capability on appearance alteration, and introduces frame-wise precise controls through techniques including DDIM inversion [41, 31] and ControlNet [52]. While achieving promising motion transfer results with variations in appearance and texture, these methods rigidly adhere to the reference frame structure and layout and fail to provide variability in the motion itself, such as adapting to new positions, intensities, camera views, or quantity of subjects.
@@
Image customization of T2I models has been explored [38, 10] where a specific unique appearance is modeled and composed into novel roles and scenes. These models are trained on a small set of images that share the same concept. They are then able to produce specific output needless of complex prompt engineering, while also allowing for diversity in poses, views, lighting, etc. compared to direct stitching and editing approaches. Inspired by this idea, we present a novel video motion customization method named Customize-A-Video based on T2V diffusion model. It customizes the model using the motion learned from the reference video, enabling it to be easily adjusted to new subjects and scenes. This includes not only precise transfer but also variations in motion intensities, positions, quantity of subjects, and camera views. These variations make the output videos more dynamic and engaging, as opposed to the robotic or unnatural appearance of per-frame replication.
@@
Specifically, we utilize a common image customization technique, Low-Rank Adaptation (LoRA) [19], applied on the pre-trained T2V diffusion model [43] to capture the motion signature in the reference video. However, applying LoRA directly to the entire T2V models proves less effective in motion preservation, as spatial and temporal characteristics are intricately entangled and both spatial and temporal information are learned simultaneously. Therefore, we apply LoRA only on temporal 3D cross-frame attention layers, creating Temporal LoRA (T-LoRA), which is more effective in capturing temporal motion dynamics from the reference video. In comparison to other popular customization techniques, LoRA also offers a portable model size and requires minimal training data, as well as the simplicity of plug-and-play for easy extensibility to collaborate with additional customization modules.
@@
While LoRA works well on few-shot customization tasks through the residual module weights, a portion of spatial features still leak into it during the training on a single reference video without diverse appearances accompanying a consistent motion concept. Concurrent efforts attempt to address this challenging yet significant issue by either concentrating on video customization demanding a small dataset with the same motion [30, 46] or stopping training early and supplementing the underfitted temporal modules with direct control signals from the reference video [20]. To tackle this issue and facilitate one-shot video customization, we introduce an innovative training approach based on a new Appearance Absorber module to further decompose spatial information from motions. The key idea of this module is that it can be trained to absorb the appearance of the reference video, leaving only the motion information for the proposed Temporal LoRA to learn.
@@
We introduce a three-stage training and inference pipeline as illustrated in Fig. 2 to connect all the components we have proposed. In the first stage, we load and train the appearance absorber on unordered reference video frames to capture frame-wise spatial information, such as the subject’s appearance and the background scene. In the second stage, we load the trained appearance absorber with frozen parameters, and train the proposed Temporal LoRA on the temporal layers of the T2V model. The appearance absorber has learned to reconstruct the static frames and therefore helps the Temporal LoRA focus primarily on temporal signals, minimizing the spatial information leaked into motion customization modules. During the final inference stage, we remove the appearance absorber that encodes the appearances from the reference video, and load solely the trained Temporal LoRA. Given a text prompt containing novel subjects and scenes, our model not only accurately transfers the learned motion signature to the new appearance, but also produces diverse motions in terms of their intensities, positions, and camera views, enhancing the dynamism and engagement of the generated videos for the motion customization task.
@@
  • We introduce a Temporal LoRA module to learn the specific motion from a reference video, facilitating motion transfer with not only accuracy but also variety;
  • We propose a new Appearance Absorber module to effectively decompose the spatial information, strategically excluding it from the motion customization process;
  • Our modules feature the plug-and-play nature and can be smoothly extended to multiple downstream applications.
@@ 2 Related Work
Text-to-Video Generation Models
@@
Text-to-video (T2V) generation task generates videos from given text prompts specifying the expected appearances and motions. It has been widely explored previously using GANs [34, 24, 22] and transformers [42, 47, 50, 18, 11]. With the emergency of T2I diffusion models, T2V diffusion models become subsequently under fast development. [21] reprogram the 2D spatial attentions into 3D temporal attentions to handle the new temporal dimension. [16, 17, 40, 59, 3, 14, 43, 26, 6] insert spatio-temporal 3D convolutions and/or cross-frame attentions to regulate the output temporal consistency from the random input noise. [28, 12, 21, 26] design explicitly disentangled noise prior between key frames and residues to enforce temporal coherency. T2V models designate the generated content through text prompts, demanding significant engineering effort to prompt it to produce desired motions in details.
T2I-based Video Editing
@@
Leveraging the control signal directly from a reference video by editing it into new appearances is a popular practice to transfer the motion and has been studied by various methods. [48, 44, 55, 36, 13, 39, 27] leverage the inverse denoising process or degraded images of the reference video frames to maintain the desired motion while altering its appearance through T2I generation. [8, 9, 45, 54, 25, 29, 1] adopt controllable image generation approaches [52, 33] and extract the low-level reference signals such as their depth or edge maps to guide the generation process. [58, 4, 7, 56, 51] make use of the combination of above techniques. However, such methods fall short as they focus more on adopting novel appearances, and merely duplicate the original motion exactly but with no temporal diversity to vary in the frame structure, motion intensity, subject position and quantity etc.
Video Motion Customization
@@
Customizing a diffusion model is the task defined as adapting the original output to a new specific domain by adjusting the pre-trained model weights. It was first introduced for T2I models to personalize in spatial aspects such as identity, art style and pose etc [38, 10, 23]. Recently, the idea of customizing the motion given reference videos has also been emerging and evolving rapidly. [48, 39, 27, 55] add temporal attentions from scratch on pre-trained T2I models and finetune them on a single video. Concurrent work [30, 20] finetunes the temporal layers in place in a pre-trained T2V model. [30] trains with special tokens following [38] on multiple videos with an extra regularization set. [20] replaces the vanilla reconstruction objective with a frame residual vector loss. Instead, our method appends residual weights to the original model based on LoRA [19] and enables advanced training strategy and inference utility.
@@
[32] represents the first attempt to finetune the spatial and temporal attentions respectively for the appearance and motion of a reference video, which are however not inferred with independently. Concurrent work [46] adds specially designed adapters over the pre-trained temporal attentions conditioned on one frame to decompose pure motion from its appearance. It requires this additional image input while ours has the minimal input requirement of text prompt only. Another concurrent work [57] applies dual-path LoRAs on spatial and temporal attentions and they are trained jointly with spatial and temporal debiased losses. In contrast, our approach adopts a staged training pipeline, where our appearance absorbers can be naturally extended to non-LoRA candidates or third-party image customization modules pre-trained on images or other videos.
Figure 3: Results of one-shot motion customization. Middle-left: ModelScope is the base T2V model which controls motions via text and fails to transfer the reference motion faithfully. Top-right: Tune-A-Video relies on DDIM inverted latent as the input and thus mostly duplicates the original frame image structure. Middle-right: Video-P2P upgrades Tune-A-Video with decouple-guidance cross-attention control but still yields deterministic output. Bottom: Our methods generates motion with both accuracy as well as variety in details such as view perspective and layout. We show two variants of our generated results on bottom left and right.
@@ 3 Method
We present a novel motion customization method based on base T2V diffusion models for a single input video. We suggest learning the motion concept from the input video through a LoRA module designed for temporal layers of the T2V model. Given the challenging nature of working with a single input video, we have developed an innovative training strategy based on an appearance absorber module to disentangle spatial information from motion. An illustration of each proposed module and its connection to the existing T2V model is shown in Fig. 2.
@@ 3.1 Preliminary
A T2V diffusion model trains a 3D UNet $\epsilon_{\theta}$ to generate videos in a series denoising steps conditioned on a input text prompt. The 3D UNet usually consists of spatial self- and cross-attentions, 2D and 3D convolutions, and temporal cross-frame attentions. Given the $F$ frames $x^{1...F}$ of a video, the 3D UNet is trained by
$L_{\theta}=\mathbb{E}_{x^{1...F},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{1...F},t,\tau(y))\|]$ (1)
at every denoising step $t=T,...,0$, where $\epsilon\sim\mathcal{N}(0,1)$ is Gaussian noise, $\tau$ is the text encoder and $y$ is the input text prompt.
Low-Rank Adaption.
Low-Rank Adaption (LoRA) [19] was proposed for adapting pre-trained large language models to downstream tasks. It has also been widely developed for image customization models. LoRA applies a residue path of two low-rank matrices $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$ in attention layers, whose original weight is $W_{0}\in\mathbb{R}^{d\times k}$, $r\ll\min(d,k)$. The new forward path is
$W=W_{0}+\alpha\Delta W=W_{0}+\alpha BA,$ (2)
@@ 3.2.2 Appearance Absorbers
We enhance the appearance absorber to separate spatial signals from temporal signals within a single video. Its objective is to absorb the spatial information, including the identity, texture, scene, etc., out of the training video, in order that the reference motion can be exclusively modeled by our T-LoRA. To achieve this, we construct the absorber using a set of image customization modules including:
Spatial LoRA (S-LoRA)
We apply LoRA on only the spatial attention layers in a T2V model to adopt solely the spatial information out of the unordered video frames. LoRA modules are injected in all self-attention layers of the frames and cross-attention layers between frames and the text prompt. We call it spatial LoRA or S-LoRA to distinguish from our T-LoRA for temporal modeling.
Textual Inversion
We utilize textual inversion [10] as an alternative approach to gather spatial features from the training video. It creates learnable placeholder tokens, initialized with briefly depicting words of the video appearance, to assimilate relevant spatial information via the pre-trained text tokenizer.
These image customization modules are adept at modeling appearance signals from limited number of frames of a single video in a few-shot manner, and thus we prefer less finetuning-based customization methods such as [38] since they require a considerable amount of training and regularization data. In addition, all types of appearance absorbers can be employed individually or jointly.
@@ 3.2.3 Training and Inference Pipelines
Our motion customization pipeline consists of two training stages for appearance absorbers and T-LoRA respectively, and one final inference stage to generate output videos with novel text prompts.
First training stage
We train appearance absorber modules first. Since they are originated from T2I models, we propose to specially train them by bypassing all temporal layers in the T2V model, including temporal attention layers and 3D convolution layers in the denoising UNet. We train them with the appearance description of the ground truth caption so that they focuses on the spatial information to learn. The input images for training are composed by unordered frames from the reference video. We follow their native loss as in [19, 10] to train each type of appearance absorber. Formally, for S-LoRA
$L_{\Delta\theta_{s}}=\mathbb{E}_{x,\epsilon,t}[\|\epsilon-\epsilon_{\theta_{0}+\Delta\theta_{s}}(x_{t}^{f},t,\tau(y))\|],$ (3)
and for textual inversion
$L_{\tau}=\mathbb{E}_{x,\epsilon,t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{f},t,\tau(y))\|].$ (4)
Second training stage
We inject previously trained appearance absorbers into the T2V model and maintain their frozen state. Our T-LoRA is meanwhile injected into the temporal attention layers of the T2V model. It is trained with the given reference video and full ground truth caption consisting of both motion verbs and appearance nouns, by which the appearance absorber is also triggered to yield spatially customized content in static frames. We use the standard reconstruction loss as employed in the diffusion model [37]:
$L_{\Delta\theta_{t}}=\mathbb{E}_{x^{1...F},\epsilon,t}[\|\epsilon-(\epsilon_{\theta_{0}+\Delta\theta_{t}}(x_{t}^{1...F},t,\tau(y)))\|].$ (5)
Inference stage
During the final inference stage, solely the trained T-LoRA is loaded onto the base T2V model. Given a new text prompt depicting the learned motion with new appearance, the customized model generates videos animated by the desired motion following the standard denoising process. As a result of the customized residual weights in T-LoRA, our output video transfers the reference motion faithfully as well as with diversity in motion intensities, positions, and camera views etc.

### v3

@@ Abstract
Image customization has been extensively studied in text-to-image (T2I) diffusion models, leading to impressive outcomes and applications. With the emergence of text-to-video (T2V) diffusion models, its temporal counterpart, motion customization, has not yet been well investigated. To address the challenge of one-shot video motion customization, we propose Customize-A-Video that models the motion from a single reference video and adapts it to new subjects and scenes with both spatial and temporal varieties. It leverages low-rank adaptation (LoRA) on temporal attention layers to tailor the pre-trained T2V diffusion model for specific motion modeling. To disentangle the spatial and temporal information during training, we introduce a novel concept of appearance absorbers that detach the original appearance from the reference video prior to motion learning. The proposed modules are trained in a staged pipeline and inferred in a plug-and-play fashion, enabling easy extensions to various downstream tasks such as custom video generation and editing, video appearance customization and multiple motion combination. Our project page can be found at https://customize-a-video.github.io.
Keywords:
Video Motion Customization Text-to-Video Diffusion Models Low-Rank Adaptation
footnotetext: * Major work was done during internships at Adobe.
@@ 1 Introduction
Replicating an iconic motion in novel scenes is highly desirable for video creation. Recent large-scale diffusion-based text-to-video (T2V) generation models [45, 5] demonstrate impressive outcomes in generating imaginative videos based on text depictions. However, they struggle with precise motion control and often demand extensive prompt engineering. Another thread of work on video editing  [50, 9, 47, 4] leverages large image generative models for appearance alteration, and introduces frame-wise precise controls via DDIM inversion [42, 32] or ControlNet [54]. While achieving promising motion transfer results with variations in appearance and texture, these methods rigidly adhere to the reference frame structure and layout and fail to provide variability in the motion itself, such as new positions, intensities, camera views, or quantity of subjects.
@@
Image customization of T2I models has been widely explored [39, 10] where a specific unique appearance is modeled and composed into novel roles and scenes. These modules are trained on a small set of images that share the same concept. They are then able to reproduce the desired concept needless of complex prompt engineering, while also allowing for diversity in poses, views, lighting, etc. compared to direct stitching and editing approaches. Inspired by this, we introduce a new task of video motion customization and present a novel one-shot method named Customize-A-Video (Fig. 1) built upon T2V diffusion models. It customizes the pre-trained model with the motion learned from the reference video, enabling it to be easily adapted to new subjects and scenes. This includes not only precise transfer but also variations in motion intensities, positions, quantity of subjects, and camera views. These variations make the output videos more dynamic and engaging, as opposed to the robotic rhythm or unnatural appearance of frame-wise tampering.
Figure 1: Customize-A-Video takes as input a single reference video (top left) and transfers its motion onto new generated videos with plausible variance. (1) Transferring the dancing twirling from the lady onto Ironman with two random output variants. (2) Transferring the motion onto multiple subjects. (3) Combining multiple motion customization together, i.e., both dancing twirling and with aerial camera flight over. (4) Combining proposed motion customization and existing image customization methods ([21] in the example) to support both appearance and motion customization.
@@
Specifically, we start from utilizing a common customization technique, Low-Rank Adaptation (LoRA) [19], applied on a pre-trained T2V diffusion model [45] to capture the motion signature in the reference video. Applying LoRA directly to the entire T2V models proves less effective in motion preservation, as spatial and temporal characteristics are intricately entangled and both will be learned simultaneously. Therefore, we apply LoRA only on temporal cross-frame attention layers, creating Temporal LoRA (T-LoRA), which is more concentrated on capturing motion dynamics from the video. In comparison to other popular customization algorithms, LoRA also offers a portable model size and requires minimal training data, as well as the simplicity of plug-and-play for easy extensibility to collaborate with additional customization modules.
@@
While LoRA works well on few-shot customization tasks through the residual module weights, a portion of spatial features still leak into it when trained on a single reference video. Concurrent efforts attempt to address this challenging yet significant issue by either demanding a small dataset with diverse appearances and the consistent motion  [31, 48], or stopping training early and supplementing the underfit temporal modules with direct control signals from the reference video [20]. To tackle this issue and facilitate one-shot video customization, we introduce an innovative Appearance Absorber module to further decompose static signals from dynamics. The key idea of this module is to absorb the appearance out of the reference video, leaving only the desired motion information for the Temporal LoRA to model.
@@
We introduce a staged training and inference pipeline as illustrated in Fig. 2 to connect all the components we have proposed while keeping them independent. In the first stage, we build and train the appearance absorber on unordered reference video frames to capture frame-wise spatial information, such as the subject appearance and the background scene. In the second stage, we load the trained appearance absorber in frozen state, and construct the Temporal LoRA on the temporal layers of the T2V model to train. The appearance absorber has encoded the static frames and therefore helps the Temporal LoRA focus primarily on temporal signals, minimizing the spatial information leakage into motion customization modules. During the inference stage, we remove the appearance absorber and load solely the trained Temporal LoRA. Given a text prompt containing novel subjects and scenes, our model not only accurately transfers the learned motion signature to the new appearance, but also produces diverse motions in terms of their intensities, positions, and camera views.
@@
  • We introduce Temporal LoRA to learn the motion from a single reference video, facilitating motion transfer with not only accuracy but also variety;
  • We propose the general class of Appearance Absorbers to dedicatedly decompose the spatial information out of the reference video, effectively excluding it from the motion customization process;
  • Our modules feature the plug-and-play and staged fashion and can be smoothly extended to various downstream applications.
@@ 2 Related Work
Text-to-Video Generation Models.
@@
Text-to-video (T2V) generation task generates videos from given text prompts specifying the expected appearances and motions. It has been widely explored previously using GANs [35, 25, 23] and transformers [44, 49, 52, 18, 11]. With the boost of T2I diffusion models, T2V diffusion models become subsequently under fast development. [22] reprogram the 2D spatial attentions into 3D temporal attentions to handle the new temporal dimension. [17, 16, 41, 62, 3, 14, 45, 27, 6] insert spatio-temporal 3D convolutions and/or cross-frame attentions to regulate the output temporal consistency from the random input noise. [29, 12, 22, 27] design explicitly disentangled noise prior between key frames and residues to enforce temporal coherency. T2V models designate the generated content through text prompts, demanding significant engineering effort to prompt it to produce desired motions in details.
T2I-based Video Editing.
@@
Leveraging the control signal directly from a reference video by editing it into new appearances is an efficient practice to precisely transfer the motion and has been studied by various methods. [50, 46, 58, 37, 13, 40, 28] leverage the inverse denoising process or degradation of the reference video frames to maintain the desired motion while altering its appearance through T2I generation. [8, 9, 47, 56, 26, 30, 1] adopt controllable image generation approaches [54, 34] and extract the low-level reference signals such as their depth or edge maps to guide the generation process. [61, 4, 7, 59, 53] make use of the combination of above techniques. However, such methods fall short as they focus more on adopting novel appearances, and merely duplicate the original motion exactly but with no temporal diversity to vary in the motion intensity and velocity, subject position and quantity etc.
Video Motion Customization.
@@
Model customization is the task of adapting the original output to a new specific domain by adjusting the pre-trained model weights. It was first introduced for T2I models to personalize in spatial aspects such as identity, art style or pose [39, 10, 24]. Recently, the idea of customizing the motion given reference videos has also been emerging and evolving rapidly. [50, 40, 28, 58] add temporal attentions from scratch on pre-trained T2I models and finetune them on a single video. Concurrent work [31, 20] tunes the temporal layers in place in a pre-trained T2V model with either a regularization set or a frame residual loss to reduce the impact of training videos’ appearances. Instead, our method appends residual weights to the original model using LoRA [19] and enables ligthweight training strategy and flexible inference utility.
@@
[33] represents the first attempt to finetune the spatial and temporal attentions independently for the appearance and motion of a reference video, which have however entangled inference. Concurrent work [57] employs two parallel UNets and tune one of them with an appearance normalization loss to disentangle the motion as well as maintain the generalization ability on new appearances. Another concurrent work [48] adds adapters conditioned on one frame to decompose pure motion from its appearance, requiring additional input of an image when inference while ours asks for the minimal input of text prompt only. Concurrent work [60] applies dual-path LoRAs and trains them jointly with an appearance-debiased loss. In contrast, our approach adopts a staged training pipeline with independent tuning configurations, where our appearance absorber class can be easily extended to more candidates than LoRA, or reusing third-party modules pre-trained on in-the-wild images or videos.
Figure 2: Our Temporal LoRA, Appearance Absorbers and their training and inference processes. All noise and denoising schedules are omitted for simplicity. (1) We bypass all temporal layers in a base T2V diffusion model and apply appearance absorber such as S-LoRA or Textual Inversion on its spatial attention layers. The module is trained on unordered video frames. (2) We apply T-LoRA on all temporal attentions in the full base T2V model. The trained appearance absorber is also loaded and frozen. The module is trained on the target video. (3) During inference, only the trained T-LoRA is loaded. A new video with the customized motion is generated by a prompt describing the new appearance and the target motion.
@@ 3 Method
We present a novel motion customization method based on pre-trained T2V diffusion models for a single reference video. We suggest learning the motion concept from the reference video through a LoRA module designed for temporal layers of the T2V model. Given the challenging nature of working with a single datum, we develop a staged training strategy with an appearance absorber module to disentangle spatial information from motion. Fig. 2 shows an illustration of each proposed module and its connection to the base T2V model.
@@ 3.1 Preliminary
A text-to-video (T2V) diffusion model trains a 3D UNet to generate videos in a series denoising steps conditioned on a input text prompt. The 3D UNet usually consists of spatial self- and cross-attentions, 2D and 3D convolutions, and temporal cross-frame attentions. Given the $F$ frames $x^{1...F}$ of a video, the 3D UNet $\theta$ is trained by minimizing
$L_{\theta}=\mathbb{E}_{x^{1...F},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{1...F},t,\tau_{v}(y))\|]$ (1)
at every denoising step $t=T,...,0$. $\epsilon$ is Gaussian noise and $\epsilon_{\theta}$ is the UNet prediction, $\tau$ is the text encoder with token sets $v$, and $y$ is the text prompt.
Low-Rank Adaptation.
Low-Rank Adaptation (LoRA) [19] was proposed for adapting pre-trained large language models to downstream tasks. It has also been widely developed for image customization models. LoRA applies a residue path of two low-rank matrices $\theta_{B}\in\mathbb{R}^{d\times r},\theta_{A}\in\mathbb{R}^{r\times k}$ on an attention layer, whose original weight is $\theta_{0}\in\mathbb{R}^{d\times k}$, $r\ll\min(d,k)$. The new forward path is
$\theta=\theta_{0}+\alpha\Delta\theta=\theta_{0}+\alpha\theta_{B}\theta_{A}.$ (2)
@@ 3.2 Customize-A-Video
We proposed two critical modules to customize the pre-trained T2V model for a single reference video. Temporal LoRA is introduced to learn motion from a reference video, whereas Appearance Absorbers are crafted to improve the separation of spatial and temporal information within the single reference video.
@@ 3.2.2 Appearance Absorbers
To separate spatial signals from temporal signals within a single video, we abstract the general class of Appearance Absorbers. Its objective is to absorb the spatial information, including the identity, texture, scene, etc., out of the training video, such that the reference motion can be exclusively modeled by our T-LoRA. To achieve this, we construct the absorbers leveraging a set of image customization modules, and use them in an inverse manner compared to their original role:
[…]
Our appearance absorbers can be built upon including but not limited to:
  • Spatial LoRA. We apply LoRA on only the spatial attention layers in a T2V model to adopt solely the spatial information out of the video frames. LoRA modules are injected in all self-attention layers of the frames and cross-attention layers between frames and the text prompt. We call it spatial LoRA (S-LoRA) to distinguish from our T-LoRA for temporal modeling.
  • Textual Inversion. We utilize textual inversion [10] as another approach to collect spatial features from the training video. It creates learnable placeholder tokens, initialized with briefly depicting words of the video appearance, to assimilate relevant spatial information through the text tokenizer.
These image customization modules are adept at modeling appearance signals from limited number of frames of a single video in a few-shot manner, and thus we prefer less finetuning-based customization methods such as [39] since they require a considerable amount of training and regularization data. All types of appearance absorbers can be employed individually or jointly.
@@ 3.2.3 Training and Inference Pipelines
Our motion customization pipeline in Fig. 2 consists of two training stages for appearance absorbers and T-LoRA respectively, and one inference stage to finally generate output videos with novel text prompts. Our configurable pipeline has dedicated stage for each module and is universal for extensive types of its components.
First training stage.
We train appearance absorber modules first. Since they are originated from T2I models, we propose to specially train them by bypassing all temporal layers in the T2V model, including temporal attention layers and 3D convolution layers in the denoising UNet. We train them with the appearance description $y_{S}$ cut out of the full caption so that they focus on learning the spatial information. The training images are unordered frames of the reference video. We follow their native loss as in [19, 10] to train each type of appearance absorber. Formally, for S-LoRA $\Delta\theta_{S}$:
$L_{\Delta\theta_{S}}=\mathbb{E}_{x,\epsilon,t}[\|\epsilon-\epsilon_{\theta_{0}+\Delta\theta_{S}}(x_{t}^{f},t,\tau_{v_{0}}(y_{S}))\|],$ (3)
and for textual inversion $\Delta v$:
$L_{\Delta v}=\mathbb{E}_{x,\epsilon,t}[\|\epsilon-\epsilon_{\theta_{0}}(x_{t}^{f},t,\tau_{v_{0}+\Delta v}(y_{S}))\|].$ (4)
Second training stage.
We inject above trained appearance absorbers into the T2V model and maintain their frozen state. Our T-LoRA is meanwhile injected into the temporal attention layers of the T2V model. It is trained with the reference video and full ground truth caption consisting of both motion verbs and appearance nouns, by which the appearance absorber is also triggered to yield spatially customized content in static frames. We train T-LoRA $\Delta\theta_{T}$ using the standard reconstruction loss as in diffusion models [38]:
$L_{\Delta\theta_{T}}=\mathbb{E}_{x^{1...F},\epsilon,t}[\|\epsilon-(\epsilon_{\theta^{\prime}+\Delta\theta_{T}}(x_{t}^{1...F},t,\tau_{v^{\prime}}(y)))\|]$ (5)
where $\theta^{\prime}=\theta_{0}+\Delta\theta_{S}$ and $v^{\prime}=v_{0}+\Delta v$ if respective AAs are employed.
Inference stage.
During the final inference stage, solely the trained T-LoRA is loaded onto the base T2V model. Given a new text prompt depicting the reference motion with new appearances and scenes, the customized model generates novel videos animated by the desired motion following the standard denoising process. As a result of the customized weights in T-LoRA, our output video transfers the reference motion faithfully as well as with diversity in motion intensities, positions, and camera views etc.
Figure 3: Results of one-shot motion customization. (1-left) Reference video. (2-left) ModelScope [45] fails to transfer the reference motion faithfully with only text guidance. (1-right & 2-right) Tune-A-Video [50] and Video-P2P [28] rely on DDIM inverted latent input and duplicate the original frame structure deterministically. (3) Concurrent work MotionDirector [60] also generates various output following the reference motion while there exist some appearance and motion artifacts especially for hard examples with complex or intensive movements. (4) Our methods generate motion with both accuracy and variety in details such as view perspective and frame layout. Two variants generated with random noise are shown for MotionDirector and Ours.